# GVDepth: Zero-Shot Monocular Depth Estimation for Ground Vehicles based on Probabilistic Cue Fusion

### v1

@@ 4.2 Ablation studies
Evaluation of canonical representations. In Tab. 3 we report extensive evaluation results, spanning different representations and train-test dataset combinations. Moreover, in Fig. 5 we report qualitative results accompanying the table. By carefully examining the results, we can draw the following conclusions: i) The baseline model struggles to transfer effectively to out-of-distribution data. Although baseline models are trained with diverse perspectives through geometric augmentations, they still rely heavily on environmental cues specific to the training domain ; (ii) Vertical image position cues generalize better than object size cues. Models with our novel $\text{VCT}_{\bm{\mathcal{C}}}(\cdot)$ representation generally outperform $\text{FCT}_{\bm{\mathcal{C}}}(\cdot)$ models, except in rare cases, potentially stemming from calibration errors. This is particularly evident with models trained on the Waymo dataset, which encounter difficulties when relying solely on object size cues. The probable cause is that Waymo’s cameras are positioned significantly higher, causing a variation of apparent object sizes due to the change of perspective. In contrast, vertical image position cues are more robust, implicitly encoding camera height within the representation; (iii) Probabilistic fusion improves generalization by integrating both cues. Our fusion strategy achieves the most accurate predictions by adaptively weighting the more reliable cue. For example, when object size cues are more accurate (e.g. DDAD $\rightarrow$ KITTI), the model adjusts accordingly, and the same applies for vertical position cues (e.g. Waymo $\rightarrow$ Argoverse).
KITTI Waymo DDAD
Refer to caption Refer to caption Refer to caption Refer to caption
Baseline Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Focal Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Vertical Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fusion Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 6: Error per object class. Vertical – model with $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. Focal – model with $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$. Both models are evaluated on DrivingStereo $\rightarrow$ KITTI zero-shot transfer.
Ablation of fusion strategy.
Table 4: Ablation of fusion strategies. Mean – simple mean instead of uncertainty fusion as in Eq. 4. L1 – loss without uncertainty weighting in Eq. 6. All models are trained on DrivingStereo.
Configuration DrivingStereo KITTI Waymo
@@
Mean 3.05 99.4 8.21 91.1 14.01 82.8
L1 3.09 99.4 8.02 91.4 13.23 82.9
Fusion 3.07 99.5 6.96 92.7 12.15 83.1
Generalization in specific regions. The strong quantitative generalization results of vertical position cues do not fully capture their performance. For instance, a model with $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ might only perform better on road pixels due to its implicitly encoded planarity. However, as shown in Fig. 6, this is not the case: the $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ model exhibits accurate generalization across a wide range of object classes.
Figure 7: Sensitivity analysis. All models are evaluated on DrivingStereo $\rightarrow$ KITTI. Histograms represent data distribution. a) $\mathrm{A.Rel}$ of $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ and $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$ models depending on ground plane modulation severity. Ground plane modulation is expressed as $\mathrm{A.Rel}$ of ground depth calculated from camera parameters via Eq. 3 and ground-truth depth of the road. b) Ratio of depth errors $(\mathrm{A.Rel}(\mathbf{Z_{Y}})/\mathrm{A.Rel}(\mathbf{Z_{F}})$ and uncertainties $(\mathbf{\Sigma}_{Y}/\mathbf{\Sigma}_{Z})$ depending on depth bin.
@@ 5 Conclusion and Future Work
In this work, we present GVDepth, a novel MDE model capable of accurate zero-shot transfer across varying environmental domains and camera setups. Our approach is tailored for autonomous vehicles and mobile robotics, where a fixed relation between the camera system and the ground plane introduces useful perspective geometry constraints. To leverage these constraints for zero-shot generalization, we proposed a Vertical Canonical Representation that uses known depth of the ground plane to enforce invariance to specific perspective geometry, resulting in enhanced learning and generalization. Additionally, we introduced a novel architecture that adaptively fuses depths obtained with object size and vertical image position cues, guided by dynamically estimated uncertanties.
Systematic ablation studies across five autonomous driving datasets demonstrated that our representation significantly improves generalization compared to object size cues, commonly used in existing zero-shot methods. Overall, our approach achieves generalization accuracy comparable to or exceeding state-of-the-art zero-shot models, all while training on a single dataset and requiring significantly fewer computational and data resources. In future work, we plan to extend this approach to multi-dataset and large-scale learning, moving toward a comprehensive foundation model for MDE in autonomous driving and mobile robotics.
@@
A Model architecture
Figure 5: Architecture of our decoder blocks. We adopt a standard upsampling architecture commonly used in U-Net encoder-decoder models with skip connections. The key distinction in our approach lies in incorporating the known depth of the ground plane as positional embeddings, similarly as in [47, 23]. This technique enriches the model with additional contextual information about the camera setup, enhancing its spatial awareness. $\mathcal{F}_{D}^{i}$ denotes the i-th decoder feature map, $\mathcal{F}_{E}^{i}$ the i-th encoder feature map, while $\bm{\mathcal{C}}=\{f_{y},c_{y},h,\theta\}$ represents the camera parameters. Each convolution operation in the architecture is followed by a GELU activation function [19].
@@ A Model architecture
Decoder architecture details. Our decoder, illustrated in Fig. 5, consists of four blocks that progressively upsample and fuse encoder features from a resolution of $(\frac{H}{32},\frac{W}{32})$ to $(\frac{H}{2},\frac{W}{2})$, while integrating 16-dimensional ground plane embeddings, similarly as in [47, 23]. Decoder channels dimensions are $\{756,512,256,128\}$. Our fusion module processes $(\frac{H}{2},\frac{W}{2})$ resolution feature maps $\mathbf{F}$ with two compact U-Net-like networks [37], each containing 3 downsampling and upsampling blocks, producing two final feature representations. These feature representations are processed with two convolutional blocks which predict our canonical representations $(\mathbf{C_{F}},\mathbf{C_{Y}})$ and accompanying uncertainty estimates $(\mathbf{\Sigma_{F}},\mathbf{\Sigma_{Y}})$. After uncertainty-based fusion, final depth map is bilinearly upsampled to $(H,W)$ resolution.
Figure 6: Vertical Canonical Representation regression. Visualization of regression boundaries of proposed Vertical Canonical Representation. We predict vertical image position of ground plane projection in range $[0,y_{MAX}]$. Image is virtually extended beyond bottom boundary to enable depth estimation for pixels with ground-plane projection below camera field-of-view. For stability reasons, regression is bounded to $y_{MAX}$, which corresponds to $d_{MAX}=80$.
@@ B Model Complexity
Figure 7: Resolution adaptability on KITTI dataset. $\mathrm{A.Rel}(\%)$ of UniDepth [35] and GVDepth as height resolution $H$ decreases from base resolution 462. Computational complexity measured by inference time on NVIDIA RTX A6000 is depicted with different circle radius. GVDepth results are from DrivingStereo $\rightarrow$ KITTI model. Note: GVDepth is adaptable only to resolutions used during training.
Why resolution adaptability matters. In our main experiments, we tackled the issue of resolution adaptability in current state-of-the-art (SotA) zero-shot methods. For instance, UniDepth [35] requires resizing and padding input images to a fixed resolution of 416×640 to achieve optimal accuracy. However, when applied to datasets like KITTI, this approach results in the model processing nearly twice as many pixels as necessary due to padding. This significantly increases computational complexity without adding any meaningful detail to the image.
C Dataset details
@@
D Camera setup calibration

### v2

@@ 4.2 Ablation studies
Evaluation of Canonical Representations. In Tab. 3 we report extensive evaluation results, spanning different representations and train-test dataset combinations. Moreover, in Supplementary we report qualitative results accompanying the table. By carefully examining the results, we can draw the following conclusions: i) The baseline model struggles to transfer effectively to out-of-distribution data. Without the invariance introduced by canonical representations, baseline models fail to leverage diverse perspective geometries in the training data due to the ambiguities induced by depth-camera parameters entanglement; (ii) Vertical image position cues generalize better than object size cues. Models with our novel $\text{VCT}_{\bm{\mathcal{C}}}(\cdot)$ representation generally outperform $\text{FCT}_{\bm{\mathcal{C}}}(\cdot)$ models, except in rare cases, potentially stemming from calibration errors. This is particularly evident with models trained on the Waymo dataset, which encounter difficulties when relying solely on object size cues. The probable cause is that Waymo’s cameras are positioned significantly higher, causing a variation of apparent object sizes due to the change of perspective. In contrast, vertical image position cues are more robust, implicitly encoding camera height within the representation; (iii) Probabilistic fusion improves generalization by integrating both cues. Our fusion strategy achieves the most accurate predictions by adaptively weighting the more reliable cue. For example, when object size cues are more accurate (e.g. DDAD $\rightarrow$ KITTI), the model adjusts accordingly, and the same applies for vertical position cues (e.g. Waymo $\rightarrow$ Argoverse).
Figure 5: Resolution adaptability. $\mathrm{A.Rel}(\%)$ of UniDepth [34] and GVDepth as height resolution $H$ decreases from base resolution 462. Computational complexity is depicted with different circle radius. GVDepth results are from DrivingStereo $\rightarrow$ KITTI evaluation.
Resolution Adaptability. SoTA zero-shot methods such as UniDepth [34] and Metric3D [49] require resizing and padding to training resolution to achieve optimal accuracy. However, when applied to datasets like KITTI, this approach results in the model processing nearly twice as many pixels as necessary due to padding. This significantly increases computational complexity without adding any meaningful detail to the image.
Table 4: Surround-view generalization. Vertical – model with $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. Focal – model with $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$. All models evaluated on DDAD. $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ and $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$ models trained solely on Waymo. (†): Waymo surround images added to training data. UniDepth and Metric3D evaluated with original checkpoints and resolutions.
Method Front-left Front-left Back
@@
Baseline 67.42 18.2 78.48 12.3 33.58 27.1
Focal 23.37 62.2 22.70 67.1 20.93 67.7
Vertical 19.21 71.8 17.22 73.5 22.12 66.2
Focal† 20.69 68.2 20.01 66.7 19.60 68.6
Vertical† 16.44 76.5 16.82 71.2 21.35 65.8
UniDepth†  [34] 72.28 8.9 71.40 8.2 55.51 11.1
Metric3D†  [49] 14.57 80.7 13.86 80.1 10.55 88.7
Figure 6: Error per object class. Vertical – model with $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. Focal – model with $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$. All models are evaluated on DrivingStereo $\rightarrow$ KITTI zero-shot transfer.
Surround-view Generalization. In Tab. 4 we provide results for Waymo $\to$ DDAD zero-shot transfer for surround-view cameras. Despite the increased ambiguity and perturbations in ground regions captured by side views, our $\operatorname{VCT(\cdot)}$ model outperforms $\operatorname{FCT(\cdot)}$ on side cameras, with performance increasing after Waymo side cameras are included in training data. $\operatorname{FCT(\cdot)}$ performs better on back camera, possibly due to the occlusion of bottom part of the image by ego-vehicle. For unknown reasons, UniDepth [34] failed to infer correct camera parameters for DDAD surround cameras, resulting in subpar performance.
Generalization in Specific Regions. The strong quantitative generalization results of vertical position cues do not fully capture their performance. For instance, a model with $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ might only perform better on road pixels due to its implicitly encoded planarity. However, as shown in Fig. 6, this is not the case: the $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ model exhibits accurate generalization across a wide range of object classes. The results clearly reveal major flaws in GEDepth’s [46] performance, as it struggles to generalize beyond road-related pixels, particularly for safety-critical objects.
Figure 7: Sensitivity analysis. All models are evaluated on DrivingStereo $\rightarrow$ KITTI. Histograms represent data distribution. a) $\mathrm{A.Rel}$ of $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ and $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$ models depending on ground plane modulation severity. Ground plane modulation is expressed as $\mathrm{A.Rel}$ of ground depth calculated from camera parameters via Eq. 3 and ground-truth depth of the road. b) Ratio of depth errors $(\mathrm{A.Rel}(\mathbf{Z_{Y}})/\mathrm{A.Rel}(\mathbf{Z_{F}}))$ and uncertainties $(\mathbf{\Sigma}_{Y}/\mathbf{\Sigma}_{Z})$ depending on depth bin.
@@ 5 Conclusion and Future Work
We present GVDepth, a novel MDE model designed for accurate zero-shot transfer across diverse environments and camera setups, with a focus on autonomous vehicles and mobile robotics. Leveraging the fixed camera-to-ground geometry in such platforms, we introduce a Vertical Canonical Representation that uses known ground depth to enforce perspective invariance, resulting in enhanced learning and generalization. Our architecture further adaptively fuses depth estimates from object size and vertical image position cues, guided by uncertainty.
Extensive ablations across five autonomous driving datasets show that our method generalizes significantly better than object-size-based baselines, matching or surpassing state-of-the-art zero-shot models, while using only a single training dataset and far fewer computational and data resources. Future work will explore scaling to multi-dataset training toward a foundation model for MDE in autonomous driving and mobile robotics.
@@
KITTI Waymo DDAD
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Baseline [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Focal [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Vertical [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Fusion [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 8: Qualitative ablation. Predicted depths and errors for DrivingStereo $\rightarrow$ {KITI, Waymo, DDAD} models. Baseline – standard depth regression with geometric augmentations. Focal – depth regression via Focal Canonical Transform - $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$. This is equivalent to Metric3D[49], but trained on our setup. Vertical – depth regression via Vertical Canonical Transform - $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. Fusion – depth regression with uncertainty-based fusion model.
F Qualitative Results
Qualitative results for different model configurations are shown in Fig. 8. All models demonstrate comparable capabilities in reconstructing semantic objects, effectively capturing object boundaries and overall scene layout. However, error maps reveal that the Baseline model struggles to handle the domain gap introduced by varying camera parameters in the zero-shot transfer setting, leading to reduced accuracy. Both the Focal and Vertical models perform reasonably well across most image regions, though they exhibit noticeable modeling errors in certain areas. In contrast, the Fusion model attains the highest accuracy by adaptively integrating the depth predictions from both Focal and Vertical models.
G Additional Ablations with Ground Plane Methods
Additional Results. Table 5 presents a more detailed ablation study of GEDepth [46], incorporating results from additional training datasets. GVDepth consistently outperforms GEDepth across multiple training-test dataset combinations. We do not report results on Argoverse and DDAD datasets, as GEDepth failed to achieve satisfactory convergence. We will provide further comments on this issue below.
Furthermore, in Tab. 6, we provide results for PlaneDepth [41] and GroCo [4], which utilize the ground plane constraint in a similar manner as GVDepth. Since these methods are self-supervised, we retrained PlaneDepth with supervision, using the same backbone as our method. The results suggest that the performance gap is not due to difference in supervision or model complexity, but due to the superior out-of-distribution generalization of our approach. For GroCo [4], we report the results from the original paper, as it lacks the open-source code implementation. However, GroCo is fundamentally a self-supervised version of GEDepth [46], preserving all of its core principles.
Our main argument is that none of these methods are explicitly designed to enhance generalization or facilitate training with diverse perspective geometries. Moreover, they improve performance only for road pixels, unlike the proposed $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ that exploits the ground plane constraint for all objects. In the following sections, we provide an in-depth analysis, offering insights into the generalization and convergence challenges faced by this methods.
Table 5: Comparison with GEDepth [46]. All models are trained with equivalent setup and model complexity on KITTI, DrivingStereo and Waymo datasets. Best results are bolded, second best are underlined. In-domain evaluation results are shaded.
Testing GEDepth[46] Vertical Fusion
dataset $\mathrm{A.Rel}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{A.Rel}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{A.Rel}\downarrow$ $\mathrm{\delta}_{1}\uparrow$
KITTI
KITTI 5.68 95.4 5.70 95.5 5.67 95.7
DStereo 18.25 66.7 10.43 87.3 10.24 87.4
Waymo 22.11 69.1 13.42 78.6 14.34 78.4
Argo 19.27 52.8 14.72 64.7 10.45 84.8
DDAD 19.46 61.2 14.26 73.6 11.81 82.5
Dstereo
KITTI 11.77 74.2 7.42 92.6 6.96 92.7
DStereo 2.99 99.5 3.07 99.5 3.01 99.5
Waymo 18.88 73.3 11.72 85.5 12.15 83.1
Argo 13.19 83.0 11.41 84.7 9.91 86.9
DDAD 20.20 55.18 15.71 80.0 12.02 82.4
Waymo
KITTI 15.91 82.1 8.30 92.3 10.92 89.8
DStereo 14.21 84.4 11.40 86.8 12.78 87.1
Waymo 3.42 99.0 3.61 98.9 3.51 98.8
Argo 17.83 83.4 12.21 88.6 9.62 94.1
DDAD 19.20 66.6 11.51 84.4 13.99 86.2
Table 6: Comparison with self-supervised methods that leverage the known ground plane. Results for KITTI $\rightarrow$ DDAD zero-shot transfer.(†): Trained with supervision. (‡): Results taken from corresponding paper.
PlaneDepth [41] 30.28 12.01 0.499 37.8
PlaneDepth [41] 32.71 14.77 0.529 37.9
GroCo  [4] 42.40 15.37 - -
Vertical† 14.26 9.39 0.271 73.6
Fusion† 11.81 8.35 0.251 82.5
Detailed Analysis.
Figure 9: Different approaches for incorporation of ground plane constraint. In our approach, $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ serves as a point of disentanglement, which enables learning of Vertical Canonical Representation that is invariant to specific perspective geometry. Linear combination used in GEDepth and GroCo does not hold such properties, leading to limited generalization.
We start with the following premises, which differentiate the proposed $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ from aforementioned approaches:
  1. (i)
    Unlike GEDepth and GroCo, which utilize the ground plane constraint only for road pixels, the proposed canonical representation enables utilization of vertical image position cue for all objects in the scene;
  2. (ii)
    Similarly, in constrast to GEDepth and GroCo, our approach is inherently modeled to enable perspective geometry disentanglement for all objects, resulting in better out-of-distribution generalization.
While our extensive evaluation confirms this, here we also provide insights into why GEDepth and GroCo fail to generalize across all objects in the scene.
Both GEDepth and GroCo formulate the problem of depth regression $\mathbf{D}\in[0,D_{max}]^{H\times W}$ as a linear combination of the known ground depth $\mathbf{D_{G}}\in[0,D_{max}]^{H\times W}$ and the learnable residual depth $\mathbf{D_{R}}\in[0,D_{max}]^{H\times W}$, weighted by the learnable “ground attention” $\mathbf{A}\in[0,1]^{H\times W}$:
$\mathbf{D}=\mathbf{A}\odot\mathbf{D_{G}}+(1-\mathbf{A})\odot\mathbf{D_{R}}.$ (9)
Through a straightforward analysis, we can notice why this formulation enhances depth estimation for road pixels only. First, any part of the scene appearing above the horizon can not benefit from the induced perspective geometry constraint, as $\mathbf{D_{G}}$ is invalid for those regions of the image. Moreover, ground attention $\mathbf{A}$ is usually estimated with the $\mathrm{Sigmoid}(\cdot)$, which has a natural tendency to converge to either 0 or 1 during optimization, leading to $\mathbf{A}$ effectively being equivalent to road segmentation, as visualized in [4, 46]. Even for rare cases where this is not the case, the resulting linear combination has an ambivalent geometric interpretation chosen arbitrarily by the model. Our proposed $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ is a simple and elegant solution for all of these issues; it enables effective utilization of the ground plane constraint for all image regions while preserving a clear geometric interpretation.
The same fundamental principles can be applied to perspective geometry disentanglement. While $\mathbf{D_{G}}$ clearly disentangles depth from camera parameters, the estimated $\mathbf{D_{R}}$ does not hold such properties, inducing generalization errors due to the ambiguity of depth and camera parameters. Since the final depth $\mathbf{D}$ is calculated as a linear combination, it is only fully disentangled where $\mathbf{A}$ is 1, which is valid only for road pixels. On the other hand, $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ is inherently designed to incorporate perspective geometry disentanglement, regardless of the specific image region.
GEDepth Convergence Issues. GEDepth optimizes $\mathbf{D}$ from Eq. 9 without additional regularization factors or external segmentation modules, claiming that the learned attention map can automatically separate ground and other regions. Unfortunately, in our experiments this occurred rather inconsistently, and the learned attention map converged to adequate ground segmentation only on the 19th training experiment. In most cases, the model resorted to estimating $\mathbf{A}=\mathbf{0}^{H\times W}$, thus ignoring the provided ground plane information and resulting in $\mathbf{D}=\mathbf{D_{R}}$.
Final Remarks. We conducted a thorough evaluation that highlights the extremely limited generalization capabilities of models that rely on ground plane constraints [41, 46, 4]. While one might assume that these issues arise from limited reproducibility or incorrect convergence, the analysis we presented in previous sections suggests that the root cause lies in inherent design choices. These models are optimized to perform well only on narrow training distribution. Even within these constraints, improvements in accuracy are confined to road pixels, which are of limited relevance for safety-critical applications. In contrast, the proposed $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ elegantly addresses all these limitations.
H Model Architecture
@@
Decoder Architecture Details. Our decoder is designed similarly as in Metric3D [49], consisting of four blocks that progressively upsample and fuse encoder features from a resolution of $(\frac{H}{32},\frac{W}{32})$ to $(\frac{H}{2},\frac{W}{2})$. Decoder channels dimensions are $\{756,512,256,128\}$. Our fusion module processes $(\frac{H}{2},\frac{W}{2})$ resolution feature maps $\mathbf{F}$ with two compact U-Net-like networks [36], each containing 3 downsampling and upsampling blocks, producing two final feature representations. These feature representations are processed with two convolutional blocks which predict our canonical representations $(\mathbf{C_{F}},\mathbf{C_{Y}})$ and accompanying uncertainty estimates $(\mathbf{\Sigma_{F}},\mathbf{\Sigma_{Y}})$. After uncertainty-based fusion, final depth map is bilinearly upsampled to $(H,W)$ resolution.
Figure 10: Vertical Canonical Representation regression. Visualization of regression boundaries of proposed Vertical Canonical Representation. We predict vertical image position of ground plane projection in range $0,y_{MAX}]$. Image is virtually extended beyond bottom boundary to enable depth estimation for pixels with ground-plane projection below camera field-of-view. For stability reasons, regression is bounded to $y_{MAX}$, which corresponds to $d_{MAX}=80$.
@@ I Model Complexity
J Ablation of Fusion Strategy.
Table 8: Ablation of fusion strategies. Mean – fusion via mean of $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$ and $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. L1 – Uncertainty loss without aleatoric uncertainty weighting in $\mathcal{L}_{unc}$. All models are trained on DrivingStereo.
Configuration DrivingStereo KITTI Waymo
Mean 3.05 99.4 8.21 91.1 14.01 82.8
L1 3.09 99.4 8.02 91.4 13.23 82.9
Fusion 3.07 99.5 6.96 92.7 12.15 83.1
K Dataset Details
@@
L Camera setup calibration
@@
M Additional Considerations
Multi-dataset Training. Almost all challenges in multi-dataset training for MDE arise from inconsistent and ambiguous perspective geometries, making our approach inherently scalable due to the invariance introduced in canonical spaces. While we lacked the computational resources to scale the training to the level of a MDE foundation model, in Tab. 3 in main text we demonstrated our method’s ability to leverage diverse perspective geometries in the training data induced by geometric augmentations, especially compared to the Baseline.
Flat Ground Assumption. Our method implicitly assumes the ideal ground plane within the $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. Unlike in GEDepth [46], which explicitly models the ground slope, we choose to grant the model greater flexibility. Motivation behind this choice is straightforward; our proposed canonical transformation is designed to assist the model in resolving ambiguities that diverse training data alone cannot, such as the entanglement of depth and camera parameters. For other adversarial perturbations, like ground plane imperfections, we do not explicitly model them. Instead, we allow the model to internally adjust values in canonical space when it detects these perturbations within highly diverse training data. Moreover, the uncertainty-based fusion provides the model with additional capabilities to adaptively weight cues based on aleatoric uncertainty, enabling the down-weighting of specific cue in highly uncertain regions.